# Enskog kinetic theory for a model of a confined quasi-two-dimensional granular fluid

### v6

@@ Appendix B Navier–Stokes collisional transfer contributions to the pressure tensor and heat flux
$P_{ij,\text{c}}^{(\text{II})}$ $=$ $-\frac{B_{1}}{d+1}n\sigma^{d}\chi\Delta^{*}\left(\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}g^{*-1}g_{x}^{*2}g_{y}^{*2}\varphi_{\text{M}}(\mathbf{c}_{1})\varphi(\mathbf{c}_{2})\right)\eta_{\text{k}}\left(\partial_{i}U_{j}+\partial_{j}U_{i}-\frac{2}{d}\delta_{ij}\nabla\cdot\mathbf{U}\right)$ (116)
@@
×$\times\left(\int\ \text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\varphi(\mathbf{c}_{1})\varphi(\mathbf{c}_{2})\left[g^{*-1}({\bf g}^{*}\cdot{\bf G}^{*})^{2}+g^{*}G^{*2}+\frac{3}{2}g^{*}({\bf g}^{*}\cdot{\bf G}^{*})+\frac{1}{4}g^{*3}\right]\right),$
@@
As in the case of the pressure tensor, the evaluation of the first term on the right hand side of Eq. (120) requires the knowledge of the first-order distribution $f^{(1)}$. By symmetry, the terms of $f^{(1)}$ contributing to $q_{i,\text{c}}^{(\text{II})}$ are $\boldsymbol{\mathcal{A}}$ and $\boldsymbol{\mathcal{B}}$. In the leading Sonine approximation, these terms are GD99 ; L05
@@
$q_{i,\text{c}}^{(\text{II})}$ $=$ $-\frac{8B_{1}}{d(d+1)(d+2)}n\sigma^{d}\chi\Delta^{*}\left(\kappa_{\text{k}}\partial_{i}T+\mu_{\text{k}}\partial_{i}n\right)\left(\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\varphi_{\text{M}}(\mathbf{c}_{1})\varphi(\mathbf{c}_{2})g^{*-1}\left[(\mathbf{g}^{*}\cdot\mathbf{S}^{*})(\mathbf{g}^{*}\cdot\mathbf{G}^{*})\right.\right.$ (122)
$\left.+g^{*2}(\mathbf{G}^{*}\cdot\mathbf{S}^{*})\right]\bigg{)}-\partial_{i}T\;\Delta^{*}\frac{m\sigma^{d+1}}{4dT}B_{2}\chi n^{2}v_{0}^{3}\left(\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\;\varphi(\mathbf{c}_{1})\frac{\partial}{\partial\mathbf{c}_{2}}\cdot(\mathbf{c}_{2}\varphi(\mathbf{c}_{2}))\right.$
@@
$=$ $-\frac{8B_{1}}{d(d+1)(d+2)}n\sigma^{d}\chi\Delta^{*}\left(\kappa_{\text{k}}\partial_{i}T+\mu_{\text{k}}\partial_{i}n\right)\left(\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\varphi_{\text{M}}(\mathbf{c}_{1})\varphi(\mathbf{c}_{2})g^{*-1}\left[(\mathbf{g}^{*}\cdot\mathbf{S}^{*})(\mathbf{g}^{*}\cdot\mathbf{G}^{*})\right.\right.$
@@ Appendix C Collision integrals
×$\times(\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g}+2\Delta)\widehat{\sigma}_{i}\alpha^{-2}f^{(0)}(\mathbf{V}_{1}^{\prime\prime})X(\mathbf{V}_{2}^{\prime\prime})$
@@
×$\times f^{(0)}(\mathbf{V}_{1})X(\mathbf{V}_{2}),$
@@
×$\times(\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g}+2\Delta)\widehat{\sigma}_{i}\Psi(\mathbf{V}_{1})\alpha^{-2}f^{(0)}(\mathbf{V}_{1}^{\prime\prime})X(\mathbf{V}_{2}^{\prime\prime})$
@@
×$\times f^{(0)}(\mathbf{V}_{1})X(\mathbf{V}_{2}).$
@@
×$\times(\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g})\widehat{\sigma}_{i}f^{(0)}(\mathbf{V}_{1})X(\mathbf{V}_{2})\left[\Psi(\mathbf{V}_{1}^{\prime})-\Psi(\mathbf{V}_{1})\right].$
@@
×$\times\int\text{d}\widehat{\boldsymbol{\sigma}}\,\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})\widehat{\sigma}_{j}\;\left[S_{j}(\mathbf{V}_{1}^{\prime})-S_{j}(\mathbf{V}_{1})\right].$
@@
$\widehat{\sigma}_{j}\;\left[S_{j}(\mathbf{V}_{1}^{\prime})-S_{j}(\mathbf{V}_{1})\right]$$\to$ $-\frac{m}{2}\Delta\left[V_{1}^{2}+2(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})^{2}-2(1+\alpha)(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})+\frac{1}{2}(1+\alpha)^{2}(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})^{2}\right.$ (143)
@@
$\times$ $\left[V_{1}^{2}+2(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})^{2}-4(1+\alpha)(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})+\frac{3}{2}(1+\alpha)^{2}(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})^{2}-3\Delta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})\right.$
@@
$-4(1+\alpha)\frac{B_{1}}{d+1}g^{-1}\left[(\mathbf{g}\cdot\mathbf{V}_{1})(\mathbf{g}\cdot\mathbf{V}_{2})+g^{2}(\mathbf{V}_{1}\cdot\mathbf{V}_{2})\right]+\frac{3}{2}(1+\alpha)^{2}B_{3}g(\mathbf{g}\cdot\mathbf{V}_{2})$
@@
$\mathcal{K}_{\kappa}^{(1)}=\frac{2^{d-1/2}d}{\sqrt{\pi}}\frac{nT^{2}}{m}\phi\chi\Delta^{*}\left[\frac{3+2d}{4}+2(1+\alpha)\left(1-\frac{3}{4}\sqrt{2\pi}\Delta^{*}\right)-3(1+\alpha)^{2}-\Delta^{*2}\right].$ (145)
@@
$\mathcal{K}_{\kappa}=2^{d}d\frac{nT^{2}}{m}\phi\chi\left\{\frac{3}{8}(1+\alpha)^{2}(1-2\alpha)+\frac{1}{\sqrt{2\pi}}\Delta^{*}\left[\frac{3+2d}{4}+2(1+\alpha)\left(1-\frac{3}{4}\sqrt{2\pi}\Delta^{*}\right)-3(1+\alpha)^{2}-\Delta^{*2}\right]\right\}.$ (146)
@@
$\times$ $\left[V_{1}^{2}+2(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})^{2}-2(1+\alpha)(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})+\frac{1}{2}(1+\alpha)^{2}(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})^{2}-3\Delta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})\right.$
@@
$\mathcal{K}_{\mu}^{(1)}=\frac{2^{d-1/2}d}{\sqrt{\pi}}\frac{nT^{2}}{m}\phi\chi\Delta^{*}\left[2\left(\Delta^{*2}+\alpha^{2}\right)-\frac{1}{2}+\frac{3}{2}\sqrt{2\pi}\alpha\Delta^{*}\right].$ (150)
@@
$\mathcal{K}_{\mu}=-2^{d}d\frac{nT^{2}}{m}\phi\chi\left\{\frac{3}{8}\alpha(1-\alpha^{2})-\frac{1}{\sqrt{2\pi}}\Delta^{*}\left[2\left(\Delta^{*2}+\alpha^{2}\right)-\frac{1}{2}+\frac{3}{2}\sqrt{2\pi}\alpha\Delta^{*}\right]\right\}.$ (151)

### v7

@@ Appendix B Navier–Stokes collisional transfer contributions to the pressure tensor and heat flux
$P_{ij,\text{c}}^{(\text{II})}$ $=$ $-\frac{B_{1}}{d+1}n\sigma^{d}\chi\Delta^{*}\left(\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}g^{*-1}g_{x}^{*2}g_{y}^{*2}\varphi_{\text{M}}(\mathbf{c}_{1})\varphi_{\text{M}}(\mathbf{c}_{2})\right)\eta_{\text{k}}\left(\partial_{i}U_{j}+\partial_{j}U_{i}-\frac{2}{d}\delta_{ij}\nabla\cdot\mathbf{U}\right)$ (116)
@@
$\times\left(\int\ \text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\varphi(\mathbf{c}_{1})\varphi(\mathbf{c}_{2})\left[g^{*-1}({\bf g}^{*}\cdot{\bf G}^{*})^{2}+g^{*}G^{*2}+\frac{3}{2}g^{*}({\bf g}^{*}\cdot{\bf G}^{*})+\frac{1}{4}g^{*3}\right]\right),$
@@
[…]
@@
[…]
$\left.+g^{*2}(\mathbf{G}^{*}\cdot\mathbf{S}^{*})\right]\bigg)-\partial_{i}T\;\Delta^{*}\frac{m\sigma^{d+1}}{4dT}B_{2}\chi n^{2}v_{0}^{3}\left(\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\;\varphi(\mathbf{c}_{1})(\mathbf{g}^{*}\cdot\mathbf{G}^{*})\frac{\partial}{\partial\mathbf{c}_{2}}\cdot(\mathbf{c}_{2}\varphi(\mathbf{c}_{2}))\right$.
@@
[…]
@@ Appendix C Collision integrals
$\times(\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g}+2\Delta)\widehat{\sigma}_{i}\alpha^{-2}f^{(0)}(\mathbf{V}_{1}^{\prime\prime})X(\mathbf{V}_{2}^{\prime\prime})$
@@
$\times f^{(0)}(\mathbf{V}_{1})X(\mathbf{V}_{2}),$
@@
$\times(\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g}+2\Delta)\widehat{\sigma}_{i}\Psi(\mathbf{V}_{1})\alpha^{-2}f^{(0)}(\mathbf{V}_{1}^{\prime\prime})X(\mathbf{V}_{2}^{\prime\prime})$
@@
$\times f^{(0)}(\mathbf{V}_{1})X(\mathbf{V}_{2}).$
@@
$\times(\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g})\widehat{\sigma}_{i}f^{(0)}(\mathbf{V}_{1})X(\mathbf{V}_{2})\left[\Psi(\mathbf{V}_{1}^{\prime})-\Psi(\mathbf{V}_{1})\right].$
@@
$\times\int\text{d}\widehat{\boldsymbol{\sigma}}\,\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})\widehat{\sigma}_{j}\;\left[S_{j}(\mathbf{V}_{1}^{\prime})-S_{j}(\mathbf{V}_{1})\right].$
@@
$\widehat{\sigma}_{j}\;\left[S_{j}(\mathbf{V}_{1}^{\prime})-S_{j}(\mathbf{V}_{1})\right]$ $\to$ $-\frac{m}{2}\Delta\left[V_{1}^{2}+2(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})^{2}-3(1+\alpha)(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})+\frac{3}{4}(1+\alpha)^{2}(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})^{2}\right$. (143)
@@
$\times$ $\left[V_{1}^{2}+2(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})^{2}-6(1+\alpha)(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})+\frac{9}{4}(1+\alpha)^{2}(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})^{2}-3\Delta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})\right$.
@@
$-6(1+\alpha)\frac{B_{1}}{d+1}g^{-1}\left[(\mathbf{g}\cdot\mathbf{V}_{1})(\mathbf{g}\cdot\mathbf{V}_{2})+g^{2}(\mathbf{V}_{1}\cdot\mathbf{V}_{2})\right]+\frac{9}{4}(1+\alpha)^{2}B_{3}g(\mathbf{g}\cdot\mathbf{V}_{2})$
@@
$\mathcal{K}_{\kappa}^{(1)}=\frac{2^{d-1/2}d}{\sqrt{\pi}}\frac{nT^{2}}{m}\phi\chi\Delta^{*}\left[\frac{3}{4}+3(1+\alpha)\left(1-\frac{1}{2}\sqrt{2\pi}\Delta^{*}\right)-\frac{9}{2}(1+\alpha)^{2}-\Delta^{*2}\right].$ (145)
@@
$\mathcal{K}_{\kappa}=2^{d}d\frac{nT^{2}}{m}\phi\chi\left\{\frac{3}{8}(1+\alpha)^{2}(1-2\alpha)+\frac{1}{\sqrt{2\pi}}\Delta^{*}\left[\frac{3}{4}+3(1+\alpha)\left(1-\frac{1}{2}\sqrt{2\pi}\Delta^{*}\right)-\frac{9}{2}(1+\alpha)^{2}-\Delta^{*2}\right]\right\}.$ (146)
@@
$\times$ $\left[V_{1}^{2}+2(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})^{2}-3(1+\alpha)(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})+\frac{3}{4}(1+\alpha)^{2}(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})^{2}-3\Delta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf V}_{1})\right$.
@@
$\mathcal{K}_{\mu}^{(1)}=\frac{2^{d-1/2}d}{\sqrt{\pi}}\frac{nT^{2}}{m}\phi\chi\Delta^{*}\left[2\Delta^{*2}-3\left(\frac{1}{2}-\alpha^{2}\right)+\frac{3}{2}\sqrt{2\pi}\alpha\Delta^{*}\right].$ (150)
@@
$\mathcal{K}_{\mu}=-2^{d}d\frac{nT^{2}}{m}\phi\chi\left\{\frac{3}{8}\alpha(1-\alpha^{2})-\frac{1}{\sqrt{2\pi}}\Delta^{*}\left[2\Delta^{*2}-3\left(\frac{1}{2}-\alpha^{2}\right)+\frac{3}{2}\sqrt{2\pi}\alpha\Delta^{*}\right]\right\}.$ (151)